# Privacy in Speech Technology

### v2

@@ III-B Attack types
III-C Attack model
In the design of privacy-protecting methods, attack models simulate and quantify an attacker’s performance. In other words, safeguards designers can put themselves in the attacker’s place to better understand weaknesses and quantify risks.
Figure 3 illustrates an attack model for measuring the extent of privacy [maouche20_interspeech]. We assume that there is some private data, which is anonymized such that it can be used in a trusted task. The attacker has access to the anonymized data, which is here called “public data”. Observe that the public data is not necessarily openly available for anyone to see, but it only indicates that the data is sufficiently freely available that the attacker has access to it. The attacker has also access to some other data about speakers, found from some other source (found data), which helps in extracting private information. Again, the term “found data” is a loose term and denotes any data other than found data that is available to the attacker.
@@ III-C Attack model
III-D Attack Surfaces
We categorize scenarios according to the attack surface from where information is extracted (see fig. 1): a channel between remote (cloud) services (section III-D1), a user interface to the local (edge) device (section III-D2) or to the cloud service (section III-D3), from the local network (section III-D4), the acoustic pathway (section III-D5), or through the shared user interface of the local device (section III-D7). We consider only cases where a piece of technology is receiving or transmitting information and exclude human-to-human communication without devices. We also assume that devices and network connections are secured such that only authorized services can communicate with them. The threats we focus on are thus illustrated with numbered, solid red lines in fig. 1.
III-D1 Cloud Leak
Suppose a user accesses a remote server, known as a cloud service through a local device, known as an edge device (see fig. 4). The cloud service is thus assumed to have legitimate access to the private messages of the user. The cloud server, however, can then use that private message or bundled side information for some other purpose than that requested by the user, extract more information than anticipated, combine it with other information, or share information with a third party. For example, the cloud server of a voice assistant could inappropriately share information with an advertiser. Observe that though figs. 4 and 1 show the secondary remote service as an impersonal cloud server, this analysis includes harms done by the people and organizations that control those servers. For the current analysis, it, however, does not make a difference who is the actor in perpetuating harm or is responsible for the harm caused to the user, but the pertinent difference is the first point where private information departs from the intended flow of information (cf. section III-D3).
@@ III-D1 Cloud Leak
III-D2 False Activation
Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig. 5). A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [wang2021wake]. The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device. Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes. They might sometimes miss a wake word when it is spoken (false negative) or mistake some other unrelated sound as a wake word (false positive). While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy. In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [liptak2017alexa, moscaritolo2018alexa].
@@ III-D3 Cloud Access
An example of this threat is medical records, which can be collected from patients to a centralized database. A researcher with authorization to access the database can then potentially extract private information beyond the expected. [Teixeira2018]
@@
III-D4 WASN Authentication
The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [das2021enhancement, zarazaga2020acoustic, sigg2020provable]. Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig. 7[Nelus2022, zarazaga2022thesis]. A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [zarazaga2022thesis, koppelmann22_interspeech]. Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup. To protect privacy, it is then necessary to determine which devices reside in the same acoustic space. Devices in a different room can belong to the same company or family and be connected to the same network, but they are still outside the sphere of the current discussion. Without proper authorization mechanisms, devices outside the room could then gain access to private speech inside the room.
@@ III-D4 WASN Authentication
III-D5 Speech Interface Leak
Figure 8 illustrates a scenario where a device or person overhears a discussion between a user, Alice, and a local device. The defining property is that the leak is in the acoustic pathway, necessitating that the eavesdropper is physically present in the same acoustic space as the private communication. The difference between these leaks and False Activation (sections III-D2 and 5) is that here, the eavesdropper is collecting data inappropriately, whereas in False Activation, an incorrect functionality is activated.
@@ III-D5 Speech Interface Leak
III-D6 Discussion Leak
Figure 9 illustrates a discussion between two persons overheard by a local device and, like False Activation, triggers unintended functionalities in the local device. The main difference to a Speech Interface Leak (sections III-D5 and 8) is that the type of speech is different. Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character. This is potentially beneficial in the sense that the speaking style of human-to-human interaction could be recognized, preventing unintended triggering of functionalities.
@@ IV Protections
The main avenues for protecting privacy in speech technology are: 1. Information isolation, where private information is filtered or otherwise removed from a transmitted (or stored) channel of speech information (section IV-A), 2. Secure processing, where access to private information is prevented with methods of cryptography (section IV-B), 3. Privacy-preserving architectures, where private information is not sent to the cloud, but processed either only locally (on the edge), or in a distributed network of (nearby) devices (section IV-A3), 4. Acoustic interventions, where reproduction of audio is changed such that private information is not revealed to eavesdroppers (section IV-D), and 5. Improving performance, where the intended functionality of a system is improved such that processing errors leak less private information (section IV-E).
IV-A Information isolation
By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information. Such isolation can be enforced through two main approaches; 1. we can reduce the capacity of the channel such that it can represent only the target information and has no room for other private information, which is known as an information bottleneck [tishby2000information] (see section IV-A1), or, 2. we can use adversarial training of the model, where the model is trained such that an attacker cannot extract a specific type of private information from the channel (see section IV-A2). Such isolation can be implemented for any category of information. More generally, a system can include multiple channels of information, each designed to carry a distinct category of information, in an approach known as disentanglement. This allows applications to cherry-pick the private information that should be transmitted case by case (see section IV-A3).
@@ IV-A Information isolation
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption OutputLabelLocal device Remote service !
@@
Figure 11: Training of a privacy-preserving speech analysis method with the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption Output--!$+$
@@
Figure 12: Training of a privacy-preserving an autoencoder structure as an example of the information bottleneck principle. The point of attack is indicated by the red arrow and exclamation mark.
IV-A1 Information Bottleneck
@@
The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed.222 A concept related to information bottlenecks is the privacy funnel [makhdoumi2014information]. Where the original definition of the information bottleneck [tishby2000information] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility. The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases. A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task. The overall approach is generic and can be applied for example in cases where 1. the decoder resynthesizes the signal with some attributes anonymized [wang2019vector] or 2. information extraction scenarios such as classification [tishby2000information], where the decoder can often be trivial.
@@
To distill only the private message and discard any side-information, speech information is passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [Tishby2015, makhdoumi2014information] (see fig. 11). This approach thus provides protection against an attacker who has access to the output of the bottleneck. The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded.
@@
A frequently used approach to implementing a bottleneck is an autoencoder structure, consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [malekzadeh2018replacement, perero2022x, lu2013speech] (see fig. 12). However, to be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in their waveform with respect to some distance function. Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original. One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [neumann2019improving, chorowski2019unsupervised].
@@
Information content at the bottleneck can be reduced by either reducing the signal rank (viz. dimensionality reduction, e.g., [perero2022x]) or by quantizing and coding the signal at a low bitrate (e.g., [wang2019vector, williams2021revisiting]). The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section V). Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel. Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight. As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
@@
A central advantage of the information bottleneck approach is that it gives a provable level of privacy. If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$. This applies to single vectors of target information passed through a bottleneck with full accuracy. Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [kairouz2015composition]. This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [nautsch20_interspeech, vali2024porcupine].
@@
Information bottlenecks, or more accurately, privacy funnels, are theoretically well-suited for privacy-preserving applications. They have also been used in other areas, such as privacy-aware transmission of images [sun2024privacyaware] and face recognition [razeghi2024deep]. However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.
@@
InferenceTraining User Encoder Trusted Task Trusted Loss Refer to caption Output Threat Task Threat Loss !$+$--
@@
IV-A2 Adversarial Approach
An alternative to reducing the size of the information bottleneck is to use an adversarial model during training, such that side-information in the bottleneck is minimized [noe21_interspeech, srivastava19_interspeech, nelus2018itg]. Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the private message and side-information. The threat task is independently optimized to extract private side information from the bottleneck output. While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^3] Accuracy refers here to some similarity measure between the message and a corrupt version of the message. of the private message and to minimize the accuracy of private side information.
Adversarial training thus combines maximizing and minimizing the accuracy of private side information. This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information. Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [nelus19b_interspeech]. This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@ IV-A2 Adversarial Approach
The principal issue with adversarial training is that it does not give any formal assurances of privacy (cf. section V-A). The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the chosen category of private information. We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
User EncoderBottleneckDecoderRefer to caption OutputAnonymization!
Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category. The point of attack is indicated by the red arrow and exclamation mark.
IV-A3 Disentanglement
With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information. For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [zhou2021vaw]. Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig. 14). Note that provable privacy requires that such channels are statistically independent of each other. That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g. [aloufi2020privacy, stoidis21_interspeech, zhou2021vaw, williams2022learning].
An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other. For example, a model can be trained to match specific channels with labels in the dataset (e.g. [stoidis21_interspeech, zhou2021vaw]). Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e. the length of time over which data is integrated (e.g. [Niekerk2020, neumann2019improving, lotfidereshgi2022cognitive]). Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds. By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.

### v3

@@ IV-C Attack model
IV-D Attack Surfaces
We categorize scenarios according to the attack surface from where information is extracted (see fig. 1): a channel between remote (cloud) services (section IV-D1), a user interface to the local (edge) device (section IV-D2) or to the cloud service (section IV-D3), from the local network (section IV-D4), the acoustic pathway (section IV-D5), or through the shared user interface of the local device (section IV-D7). We consider only cases where a piece of technology is receiving or transmitting information and exclude human-to-human communication without devices. We also assume that devices and network connections are secured such that only authorized services can communicate with them. The threats we focus on are thus illustrated with numbered, solid red lines in fig. 1.
IV-D1 Cloud leak
Suppose a user accesses a remote server, known as a cloud service, through a local device, known as an edge device (see fig. 4). The cloud service is thus assumed to have legitimate access to the target information of the user. The cloud server, however, can then use that target information or bundled side information for some other purpose than that requested by the user, extract more information than anticipated, combine it with other information, or share information with a third party. For example, the cloud server of a voice assistant could inappropriately share information with an advertiser.
Another example is data collection for training of deep neural networks. A majority of advanced speech processing today uses machine learning, which has to be trained using large databases of speech. The best quality data correspond closely to scenarios where the services are used. Recording users’ interactions with their devices is then attractive for training improved models since it corresponds exactly to the use case. This presents a considerable threat to privacy because such unrestricted recording could capture a wide range of private information, including all interactions with the device but potentially also any and all speech in the vicinity of the device.
Observe that though figs. 4 and 1 show the secondary remote service as an impersonal cloud server, this analysis includes harms done by the people and organizations that control those servers. For the current analysis, it, however, does not make a difference who is the actor in perpetuating harm or is responsible for the harm caused to the user, but the pertinent difference is the first point where private information departs from the intended flow of information (cf. section IV-D3).
@@ IV-D1 Cloud leak
IV-D2 False activation
Devices with speech interfaces can hear all conversations in the same acoustic space as the device and, therefore, need mechanisms to determine which utterances are intended for the speech interface (see fig. 5). A popular approach is to use a specific utterance, known as a wake word, to start all interactions with the interface [105]. The wake word is then like a rudimentary password, which prevents the interface from activating when speech is not directed to the device. Unfortunately, designing wake word detectors is non-trivial, and they will occasionally make mistakes. They might sometimes miss a wake word when it is spoken (false negative) or mistake some other unrelated sound as a wake word (false positive). While false negatives are annoying for the user when the service does not activate, false positives potentially present serious threats to privacy. In some famous cases, speech interfaces have activated from sounds on the television to buy unwanted items at the users’ behest, and users’ private conversations have been leaked to third parties [106, 107].
@@ IV-D3 Cloud access
An example of this threat is medical records, which can be collected from patients to a centralized database. A researcher with authorization to access the database can then potentially extract private information beyond the expected [108].
@@
IV-D4 WASN authentication
The audio quality of speech pickup (in terms of, for example, signal-to-noise ratio or perceptual quality), as well as the usability of voice interfaces, can be improved by using all available connected devices with microphones in the same room or acoustic space [109, 79, 110]. Such collaboration can be realized with wireless acoustic sensor networks (WASNs), where several independent devices simultaneously pick up speech, and those channels are combined to obtain a high-quality signal (see fig. 7[16, 18]. A sensitive question is, however, authorization: Which devices are allowed to share information with each other?
One approach is to assume that devices that are in the same room or acoustic space can hear the same signal [18, 72]. Presence in the same space is thus already an implicit authorization to participate in a joint signal pickup. To protect privacy, it is then necessary to determine which devices reside in the same acoustic space. Devices in a different room can belong to the same company or family and be connected to the same network, but they are still outside the sphere of the current discussion. Without proper authorization mechanisms, devices outside the room could then gain access to private speech inside the room.
@@ IV-D4 WASN authentication
IV-D5 Speech interface leak
Figure 8 illustrates a scenario where a device or person overhears a discussion between a user, Alice, and a local device. The defining property is that the leak is in the acoustic pathway, necessitating that the eavesdropper is physically present in the same acoustic space as the private communication. The difference between these leaks and False Activation (sections IV-D2 and 5) is that here, the eavesdropper is collecting data inappropriately, whereas in False Activation, an incorrect functionality is activated.
@@ IV-D5 Speech interface leak
IV-D6 Discussion leak
Figure 9 illustrates a discussion between two persons overheard by a local device and, like False activation, triggers unintended functionalities in the local device. The main difference to a Speech interface leak (sections IV-D5 and 8) is that the type of speech is different. Interaction with a speech interface at the current level of technology is, roughly speaking, limited to speech commands, question-answer pairs, and one-sided monologues, whereas spoken human-to-human interaction has potentially a much more dynamic character. This is potentially beneficial in the sense that the speaking style of human-to-human interaction could be recognized, preventing unintended triggering of functionalities.
@@ V Protections
The main avenues for protecting privacy in speech technology are: 1. Information isolation, where private information is filtered or otherwise removed from a transmitted (or stored) channel of speech information (section V-A), 2. Secure processing, where access to private information is prevented with methods of cryptography (section V-B), 3. Privacy-preserving architectures, where private information is not sent to the cloud, but processed either only locally (on the edge), or in a distributed network of (nearby) devices (section V-A3), 4. Acoustic interventions, where reproduction of audio is changed such that private information is not revealed to eavesdroppers (section V-D), 5. Improving utility, where the intended functionality of a system is improved such that processing errors leak less private information (section V-E) and 6. Design of provable protectionsby applying theoretical frameworks such as differential privacy (section V-F).
V-A Information isolation
By isolation of information, we refer to the design of a channel of information where the target information is isolated from other (undesired) private information. This applies at least and in particular to the attack vectors Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), where a reduction in flow of private information would reduce threats to privacy. Central concepts in information isolation include 1. information bottlenecks(or privacy funnels) , where information is compressed to a low-capacity channel such that only target information is passed through (see section V-A1) and where the neural network implementation usually takes the structure of an autoencoder [111], 2. adversarial training, where an information channel (such as an information bottleneck) is trained to contain target information to fulfill the trusted task and simultaneously trained such that a competing, attacking adversarial model fails to extract private information (threat task), (see section V-A2) and 3. disentanglement, where different attributes of speech are isolated to separate channels, as a generalization of information bottlenecks (see section V-A3). Note that these concepts are overlapping and should be seen as characterizations of algorithms and methodology, rather than alternative approaches.
@@ V-A Information isolation
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption OutputLabelLocal device Remote service !!
@@
Figure 11: Training of a privacy-preserving speech analysis method with the information bottleneck principle. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
InferenceTraining User EncoderBottleneckDecoder Loss Refer to caption Output-!++!
@@
Figure 12: Training of a privacy-preserving an autoencoder structure as an example of the information bottleneck principle. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
V-A1 Information bottleneck
@@
The purpose of an information bottleneck is to use an encoder to distill the information needed to perform a downstream (trusted) task into a bottleneck, such that all other (private) information is removed. A decoder can then be used to decipher the information in the bottleneck into a format usable for the downstream task. The overall approach is generic and can be applied, for example, in cases where 1. the decoder resynthesizes the signal with some attributes anonymized [112] or 2. information extraction scenarios such as classification [113], where the decoder can often be trivial.
A concept related to information bottlenecks is the privacy funnel [114]. Where the original definition of the information bottleneck [113] aims to maximize the utility of a channel with a fixed rate, the privacy funnel aims to minimize the information rate with a fixed utility. The information bottleneck is, however, much more widely known in the community and often informally used as an umbrella term to cover both cases.
@@
To distill only the target information and discard any side-information, speech information can be passed through a bottleneck so tight, that only the legitimate message at full accuracy can pass through [115, 114] (see fig. 11). This approach thus provides protection against an attacker who has access to the output of the bottleneck. The challenges are to design a training methodology and model structure with a sufficiently tight bottleneck so that the legitimate message is accurately retained while private side information is discarded. Observe that a generic implementation of an information bottleneck does not guarantee privacy as side-information can sneak through the bottleneck alongside the target information [116]. Instead, to be privacy preserving, a channel must be designed or optimized for privacy, warranting the term privacy funnel.
@@
A frequently used approach to implementing a bottleneck is an autoencoder structure [111], consisting of an encoder, bottleneck, and a decoder, where the objective is to reconstruct the input signal from the bottleneck output [78, 49, 117] (see fig. 12). To be privacy-preserving, the bottleneck should be sufficiently tight that the original speech signal cannot be perfectly reconstructed to resemble the original in its waveform with respect to some distance function. Designing a loss function is then challenging when the bottleneck is so tight that the decoded signal is very far from the original. One solution akin to the machine learning method known as representation learning is to feed the output of the autoencoder again to the encoder and compare the bottleneck features [118, 119].
@@
Information content at the bottleneck can be reduced by either reducing the signal rank (viz. dimensionality reduction, e.g., [49]) or by quantizing and coding the signal at a low bitrate (e.g., [112, 120]). The trade-off between the accuracy of trusted task and bottleneck entropy then determines the extent of privacy (see section VI). Observe that dimensionality reduction is, in general, not alone sufficient because if the channel’s floating- (or fixed)-point representation has $K$ dimensions each with $B$ bits, then a sufficiently advanced encoder can pack $KB$ bits of information to that channel. Since typical representations have $B=32$ or $B=64$, which can encode up to $2^{B}$ unique speaker identities, we see that the bottleneck is not particularly tight. As a general rule, therefore, bottlenecks should always be quantized to the lowest bitrate sufficient to represent the target information.
@@
A central advantage of the information bottleneck approach is that it can give a provable level of privacy. If the bitrate of the bottleneck is $B_{b}$ and that of the target information is $B_{t}$, then the privacy leak is at most $B_{leak}\leq B_{t}-B_{b}$. This applies to single vectors of target information passed through a bottleneck with full accuracy. Streaming data through the bottleneck, however, over time, leaks a flow of private information that increases in size without bound [121, 122]. This approach also addresses only the average disclosure of private information but does not give guarantees of worst-case disclosure [87, 123].
@@
Information bottlenecks, or more accurately, privacy funnels, are thus theoretically well-suited for privacy-preserving applications. They have also been used in other areas, such as privacy-aware transmission of images [124] and face recognition [125]. However, the above-mentioned applications in speech either do not specifically apply it for privacy or use information bottlenecks optimized for privacy with additional components.
@@
InferenceTraining User Encoder Trusted Task Trusted Loss Refer to caption Output Threat Task Threat Loss !++-
@@
V-A2 Adversarial training
Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [48, 64, 92]. Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the target information and side-information. The threat task is independently optimized to extract private side information from the bottleneck output. While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^2] Accuracy refers here to some similarity measure between the message and a corrupt version of the message. of the target information and to minimize the accuracy of private side information.
Adversarial training thus combines maximizing and minimizing the accuracy of private side information. This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information. Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [47]. This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@ V-A2 Adversarial training
The principal issue with adversarial training is that theoretical frameworks for certifying performance (e.g. [126]) are not yet widely accepted and there are thus no formal assurances of privacy (cf. section VI-A). The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the selected category of private information. We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
User EncoderBottleneckDecoderRefer to caption OutputAnonymization!!
Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
V-A3 Disentanglement
With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information. For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [127]. Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig. 14). Note that provable privacy requires that such channels are statistically independent of each other. That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g. [128, 93, 127, 20].
The privacy-preserving performance of a disentangling system can be readily formulated, provided that channels $k$ are uncorrelated. If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [121]. Per-channel disclosures are thus additive such that the design and performance analysis of the privacy-preserving systems is straightforward.
An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other. For example, a model can be trained to match specific channels with labels in the dataset (e.g. [93, 127]). Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e. the length of time over which data is integrated (e.g. [129, 118, 130]). Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds. By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.